OpenStudy (anonymous):

Find the exact solution and a two-decimal-place approximation exponential equation and special change of base formula: (1/3)^x = 100

sam (.sam.):

Natural log both sides

OpenStudy (anonymous):

or log base ten both sides...

sam (.sam.):

\[\ln(\frac{1}{3})^{x}=\ln(100)\]

sam (.sam.):

\[xln(1/3)=\ln100\] \[x=\frac{\ln100}{\ln(1/3)}\]

OpenStudy (anonymous):

thank you.

sam (.sam.):

yw

OpenStudy (anonymous):

Thank you also, dpalnc.

OpenStudy (anonymous):

yw